# Field-based Molecule Generation

### v1

@@ 3.3 Noise schedules
The cosine schedule for the forward process variance $\sigma_{t}^{2}$ has been shown to increase the negative log-likelihood and image quality (Nichol & Dhariwal, 2021). The effect is that data points change at a slower pace between consecutive time steps compared to the linear scheduler proposed by Ho et al. (2020a). To further emphasize the stability of generated molecules, Vignac et al. (2023) and Peng et al. (2023b) employed separate noise schedulers for the bonds and atom types, using a cosine scheduler with an additional parameter, and a new, sigmoid scheduler, respectively. Interestingly, these methods reported similar improvements while using reverse noise process order, where the bond type noising process is faster than atoms’ in Vignac et al. (2023), and vice-versa in Peng et al. (2023b).
We achieved better results using the approach of Vignac et al. (2023), and therefore employ a model that first generates the bonds (molecular structure), and then fills in the atoms. Their proposed noise scheduler is:
@@
where $\alpha_{t}$ determines the signal amplitude for a given time step $t$ in Equation 4, and we set $\nu$ to $1$ and $1.5$ for atom and bond channels, respectively. This induces a tendency to generate bonds earlier than atoms. In Figure 3, $\alpha_{t}$ values are plotted for atom and bond fields, respectively, along with an illustration of how bond RBF components $\mathbf{u}_{b}$ are generated before atoms $\mathbf{u}_{a}$, over time steps $t\in[0,1]$, for one sample generated by FMG (please refer to Appendix G for details on how this figure is created).
@@ 3.4 Classifier free guidance
Classification guidance in denoising diffusion models was introduced as a technique to control the tradeoff between the sample diversity and quality (Dhariwal & Nichol, 2021). This method requires training a separate classifier for a feature or label $\mathbf{y}$, $p_{\phi}(\mathbf{y}|\mathbf{u}_{t})$, on corrupted data points $\mathbf{u}_{t}$ at arbitrary noise scales $t$. The gradients of the log probability $\nabla_{\mathbf{u}}\log p_{\phi}(\mathbf{y}|\mathbf{u}_{t})$ are then used to guide the data points $\mathbf{u}_{t}$ in the direction of label $\mathbf{y}$ region. Following this work, Ho & Salimans (2022) empirically showed that using an approximation of the implicit classifier $p(\mathbf{y}|\mathbf{u}_{t})\propto p(\mathbf{u}_{t}|\mathbf{y})/p(\mathbf{u}_{t})$, instead of a separately trained classifier $p_{\phi}(\mathbf{y}|\mathbf{u}_{t})$ achieves similar results. Concretely, given the conditional and unconditional noise estimators $\boldsymbol{\epsilon}_{\theta}(\mathbf{u}_{t};\mathbf{y})$ and $\boldsymbol{\epsilon}_{\theta}(\mathbf{u}_{t})$, they use the following linear interpolation during sampling:
$\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{u}_{t};\mathbf{y})=(1+\beta)\boldsymbol{\epsilon}_{\theta}(\mathbf{u}_{t};\mathbf{y})-\beta\boldsymbol{\epsilon}_{\theta}(\mathbf{u}_{t}),$ (12)
where we note that $\boldsymbol{\epsilon}_{\theta}(\mathbf{u}_{t};\mathbf{y})$ and $\boldsymbol{\epsilon}_{\theta}(\mathbf{u}_{t})$ are up to a constant equal to the conditional and marginal score estimates $\nabla_{\mathbf{u}}\log p_{\theta}(\mathbf{u}_{t}|\mathbf{y})$ and $\nabla_{\mathbf{u}}\log p_{\theta}(\mathbf{u}_{t})$.
We use classification-free guidance to control the generation on the number of atoms. As expected, we observe improvements in all metrics and we use this, as it provides a better comparison to point-cloud molecule generation methods, which always require pre-specifying the number of atoms (e.g. sampled according to training data statistics).
3.5 Architecture choice and rotation invariance
We parametrize the denoiser $\boldsymbol{\epsilon}_{\theta}$ with a three-dimensional U-Net architecture, which consists of ResNet blocks, scaling layers, and attention layers. The choice is well-suited for translational invariance, as convolutions are in principle invariant to translations. However, convolutions are not rotationally invariant.
@@ 3.6 Extracting molecular graphs
To extract a molecular graph from the generated fields $\mathbf{u}=(\mathbf{u}_{\texttt{H}},\mathbf{u}_{\texttt{C}},\ldots,\mathbf{u}_{b_{1}},\mathbf{u}_{b_{2}},\mathbf{u}_{b_{3}})$, we first initialize the number of atoms $N$, their positions $\mathbf{m}_{i}$, and their types $a$, using a computationally efficient peak picking algorithm which retrieves sets $A_{a}$ of atom positions $\mathbf{m}_{i}$, for each atom type. We then optimize the mean-squared error loss between the values of the discretized generated fields $\mathbf{u}_{a}$, and their corresponding RBF function values determined by $\mathbf{m}_{i}\in A_{a}$
$\begin{split}\operatorname*{arg\,min}_{\{\mathbf{m}_{i}\}}:\sum_{a\in\mathcal{A}}&\sum_{\mathbf{x}\in\mathcal{G}}\Big{[}\mathbf{u}_{a}(\mathbf{x})-\gamma_{a}\sum_{i=1}^{N}\mathds{1}[a_{i}=a]\\ &\quad\exp\Big{(}-\frac{1}{2\sigma_{a}^{2}}||\mathbf{x}-\mathbf{m}_{i}||^{2}\Big{)}\Big{]}^{2},\end{split}$ (13)
by matching the generated atom fields $\mathbf{u}_{a}$ against the RBF mixtures induced by $\{\mathbf{m}_{i}\}$ across the grid points $\mathcal{G}$.
Similarly, given the optimized atom positions $\{\mathbf{m}_{i}\}$, we initialize the bonds (see Appendix D for details). Briefly, considering the atom positions $\mathbf{m}_{i}$ fixed, we determine the bonds by first considering atoms within lenient distances (determined based on data statistics, with large margins), and filtering for those that contain large field values $\mathbf{u}_{b}$ on any bond channel $b\in\mathcal{B}$, around the mean points $\mathbf{m}_{ij}$ determined by the (potentially) bonded atom positions $\mathbf{m}_{i}$ and $\mathbf{m}_{j}$. Denoting the initial bond sets for each channel as $B_{ij,b}$, we optimize the RBF component $\gamma_{ij}$ values, which have $B_{ij,b}\neq\varnothing$:
$\begin{split}\operatorname*{arg\,min}_{\gamma_{ij}}\sum_{b\in\mathcal{B}}&\sum_{\mathbf{x}\in\mathcal{G}}\Big{(}\mathbf{u}_{b}(\mathbf{x})-\sum_{(i,j)}\gamma_{ij}\mathds{1}[b_{ij}=b]\\ &\exp\Big{(}-\frac{1}{2\sigma_{b}^{2}}||\mathbf{x}-\mathbf{m}_{ij}||^{2}\Big{)}\Big{)}^{2},\end{split}$ (14)
where the $\gamma_{ij}$ values corresponding to incorrect bonds should optimally be close to 00.
Figure 4 illustrates a bond that was initially extracted based on $B_{ij,b}\neq\varnothing$, but removed after optimizing and finding the low $\gamma_{ij}$ value. Please refer to Appendix D for a detailed explanation of the molecular graph extraction procedure.
Figure 4: The bond marked with “x” on the left is removed since its optimized $\gamma_{ij}$ value (right bar plot) is significantly lower than the rest. The illustration shows the bond type 1 values of a noisy field $\tilde{\mathbf{u}}_{b_{1}}=\mathbf{u}_{b_{1}}+\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}$ is a tensor of the same dimension as $\mathbf{u}_{b_{1}}$, containing samples of independent uniform noise $\mathcal{U}(0,0.1)$. The bond is initially considered due to our very inclusive initial bond extraction procedure applied to a noisy field. The bond field values $\mathbf{u}_{b_{1}}$ above a threshold of $0.2$ are shown and encoded according to the color bar.
@@ 4 Evaluation
We use atom and molecule neutrality, validity, and novelty, as defined in (Guan et al., 2023). We call the stability metric introduced in (Guan et al., 2023) neutrality, to clarify and distinguish from other work that considered stable atoms even those without a formal charge of 0 (Vignac et al., 2023). Atom neutrality is the percentage of atoms that have a formal charge of 0, while molecule neutrality is the percentage of molecules having all atoms with 0 charge. Although in practice generating H atoms explicitly can be argued to be redundant, we emphasize that this still serves as one of the best benchmarking tools. Unlike validity, this metric cannot be trivially increased by e.g., extracting only covalent bond type 1 between atoms, or other such trivial solutions that may increase one metric at the expense of other distribution properties (Guan et al., 2023).
Note how the neutrality metric is strictly appropriate only when generating bond types =$\mathcal{B}=\{b_{1},b_{2},b_{3}\}$. A few methods, such as the model from (Vignac et al., 2023) and (Peng et al., 2023a) opted for explicitly generating aromatic rings, which creates ambiguity in the neutrality and validity metrics we have used. Please refer to Appendix I for details about this problem.
@@
We compute atom TVa and bond TVb total variation between the datasets and generated distributions, and report the sum of the distances over all atom and bond types, respectively.
Next, we compute the average maximum similarity between all generated molecules and the test dataset. We compute the Morgan fingerprints (Zhong & Guan, 2023) of all the generated and test dataset molecular graphs and determine the maximum Tanimoto Similarity between each generated molecule and the test set, which we denote by MST.
@@
To assess the quality of the generated 3D molecular conformations, we employ the Wasserstein metrics introduced in (Vignac et al., 2023), and compute bond angle (BA) and bond length (BL) W1 distance between generated and data distributions.
Please refer to Appendix H for more details about the performance metrics.
@@ 5.1 QM9
QM9 (Ramakrishnan et al., 2014) is a small molecule dataset, containing 134k molecules of at most 29 atoms, including hydrogen. All molecules in the data are described by their atom positions and computed geometric, energetic, electronic, and thermodynamic properties. We use the same splits defined in (Guan et al., 2023), where we split the data into 100k, 18k, and 10k molecules for training, validation, and testing, respectively.
Out of the 100k training molecules, we use 98522 molecules. The molecules were selected such that the 95% percentile of the RBF components corresponding to each atom $\mathbf{m}_{i}\pm 2\sigma$ would fit within the selected field evaluated positions $\mathcal{G}$.
In Table 1, we compare our model to a wide variety of previous methods. We achieve state-of-the-art performance in terms of molecule and atom neutrality, indicating that our method captures the basic chemical properties very well.
In Table 2, the maximum test similarity and the bond distribution are competitive to the point-cloud approaches, while MiDi performs very well on TVa. Since we lack the strong enforcement of the atom number for each generated molecule, and only use classification-free guidance to stochastically guide the model, we further analyze how well the generated molecules adhere to the number of atoms we condition the model on in Appendix F. Our generative method likely trades this property for higher molecular stability, by removing or replacing the RBF components in the atom fields $a$, such that the atoms’ formal charges are 0. We note how this is not possible for current point-cloud methods, as the number of atoms in each molecule is strictly enforced, and these methods can only change the atom types and locations that are initially randomly assigned, but not the number of atoms in the generated molecules.
[…]
Table 1: Basic molecule properties for molecular graphs generated by models trained on the QM9 dataset. Results marked with (*) use deterministic, post-processing techniques to determine the bonds or assign formal charges to their atoms when computing the metrics. Results marked with ($\dagger$) have been re-tested by us, and those marked with ()($\ddagger$)( ‡ ) generate an additional, explicit aromatic bond channel. We compute the mean and standard deviation of the metrics for three different runs, each generating 10k samples. All properties are expressed in percentages.

### v2

@@ 4.3 Noise schedules
To further emphasize the stability of generated molecules, Vignac et al. (2023b) and Peng et al. (2023) employed separate noise schedulers for the bonds, atom types, and their positions using a modified cosine and a sigmoid scheduler, respectively. We also observed increased performance using a similar approach to Vignac et al. (2023b), which adds a parameter $\nu$ to the original cosine schedule from (Nichol & Dhariwal, 2021):
@@
where $\alpha_{t}$ determines the signal amplitude for each time $t$ in Equation 4, and we set $\nu$ to $1$ and $1.5$ for atom and bond channels, respectively. The schedule $\alpha_{t}$ and its effect on generation can be seen in Figure 2, where bonds are denoised faster than atoms, increasing the generated molecules’ stability.
@@ 4.4 Classifier free guidance
Ho & Salimans (2022) empirically showed that using an approximation of the implicit classifier $p(\mathbf{y}|\mathbf{u}_{t})\propto p(\mathbf{u}_{t}|\mathbf{y})/p(\mathbf{u}_{t})$ achieves similar conditional generation performance to (Dhariwal & Nichol, 2021), without needing a separately trained classifier $p_{\phi}(\mathbf{y}|\mathbf{u}_{t})$ for multiple noise levels $t$. During sampling, the conditional and unconditional noise estimators $\bm{\epsilon}_{\theta}(\mathbf{u}_{t};\mathbf{y})$ and $\bm{\epsilon}_{\theta}(\mathbf{u}_{t})$ are linearly interpolated:
$\tilde{\bm{\epsilon}}_{\theta}(\mathbf{u}_{t};\mathbf{y})=(1+\beta)\bm{\epsilon}_{\theta}(\mathbf{u}_{t};\mathbf{y})-\beta\bm{\epsilon}_{\theta}(\mathbf{u}_{t}),$ (12)
where we note that $\bm{\epsilon}_{\theta}(\mathbf{u}_{t};\mathbf{y})$ and $\bm{\epsilon}_{\theta}(\mathbf{u}_{t})$ are up to a constant equal to the conditional and marginal score estimates $\nabla_{\mathbf{u}}\log p_{\theta}(\mathbf{u}_{t}|\mathbf{y})$ and $\nabla_{\mathbf{u}}\log p_{\theta}(\mathbf{u}_{t})$.
Point-cloud approaches always strictly enforce the generated atom counts. We determine the effect of atom count conditioning using classifier-free guidance on generated molecule statistics. While not strictly necessary for our field-based formulation, we show that atom count conditioning yields molecule graph quality improvements. Further details can be found in Appendix G.
4.5 Architecture choice, rotation invariance and scaling
We parametrize the denoiser $\bm{\epsilon}_{\theta}$ with a three-dimensional U-Net architecture, which consists of ResNet blocks, scaling layers, and attention layers. CNN layers provide useful translational invariance inductive biases but lack rotation invariance.
@@ 4.5 Architecture choice, rotation invariance and scaling
Compared to previous methods that use E(3) invariant NN and can generate molecules with any rotation, our method (on optimality) will only generate molecules according to this frame of reference. NN hyperparameters can be found in Appendix D.
Our method’s scaling is dependent on the evaluated field locations $\mathcal{G}$. In Appendix J, we derive the upper limit of our method’s scaling and show that it differs substantially from empirical values, even surpassing the generation speed of (Hoogeboom et al., 2022) for the largest molecules.
@@ 4.6 Extracting molecular graphs
We briefly illustrate the molecular graph extraction method and give additional details in Appendix E. To extract a molecular graph from the generated fields $\mathbf{u}=(\mathbf{u}_{\texttt{H}},\mathbf{u}_{\texttt{C}},\ldots,\mathbf{u}_{b_{1}},\mathbf{u}_{b_{2}},\mathbf{u}_{b_{3}})$, we first initialize the number of atoms $N$, their positions $\{\mathbf{m}_{i}\}$, and their types $\{\mathbf{a}_{i}\}$, using a computationally efficient peak picking algorithm which retrieves sets $M_{a}$ of atom positions $\mathbf{m}_{i}$, for each atom type. We then optimize the mean-squared error loss between the values of the generated fields $\mathbf{u}_{a}$, and the corresponding atom RBF function values determined by $\mathbf{m}_{i}\in M_{a}$ by matching the generated atom fields $\mathbf{u}_{a}$ against the RBF mixtures induced by $\{\mathbf{m}_{i}\}$ across the grid points $\mathcal{G}$
$\operatorname*{arg\,min}_{\{\mathbf{m}_{i}\}}\sum_{a\in\mathcal{A}}\sum_{\mathbf{x}\in\mathcal{G}}\Big{[}\mathbf{u}_{a}(\mathbf{x})-\gamma_{a}\sum_{i=1}^{N}\mathds{1}[a_{i}=a]\exp\Big{(}-\frac{1}{2\sigma_{a}^{2}}||\mathbf{x}-\mathbf{m}_{i}||^{2}\Big{)}\Big{]}^{2}.$ (13)
Similarly, given the fixed, optimized atom positions $\{\mathbf{m}_{i}\}$ we initialize bonds. All atom position pairs $(\mathbf{m}_{i},\mathbf{m}_{j})$ within certain distances define our initial bond set $B$. We then consider sets $B_{ij,b}$ of values $\mathbf{u}_{b}$ around bond positions $\mathbf{m}_{ij}=(\mathbf{m}_{i}+\mathbf{m}_{j})/2$, that are above predefined thresholds, for every $(\mathbf{m}_{i},\mathbf{m}_{j})\in B$. For all $B_{ij,b}\neq\varnothing$ we assign $\gamma_{ij}$ and optimize:
$\begin{split}\operatorname*{arg\,min}_{\gamma_{ij}}\sum_{b\in\mathcal{B}}\sum_{\mathbf{x}\in\mathcal{G}}\Big{(}\mathbf{u}_{b}(\mathbf{x})-\sum_{(i,j)}\gamma_{ij}\mathds{1}[b_{ij}=b]\exp\Big{(}-\frac{1}{2\sigma_{b}^{2}}||\mathbf{x}-\mathbf{m}_{ij}||^{2}\Big{)}\Big{)}^{2},\end{split}$ (14)
where the small RBF component values $\gamma_{ij}$ corresponding to erroneously detected bonds are removed.
@@ 5 Evaluation
We briefly present the benchmarking metrics and refer to Appendix H for details and definitions.
@@
We use atom and molecule neutrality and validity. Neutrality is the same metric defined in (Hoogeboom et al., 2022) as ”stability”, but, since other work considered stable atoms even those without a formal charge of 0 (Vignac et al., 2023b), we denote it neutrality, to clarify and distinguish from their definitions
Atom neutrality is the percentage of atoms with a formal charge of 0, while we consider molecules as neutral if all their atoms are neutral (note how the definition slightly differs from the chemical one).
Unlike validity, neutrality cannot be trivially increased by e.g., generating only covalent type 1 bonds, or other trivial solutions that may increase this metric at the expense of other distribution properties (Hoogeboom et al., 2022). We treat neutrality like any other distribution property, and we aim to match the generated neutral molecule fraction of the data.
@@
[…]
Next, we report the average maximum similarity between the test dataset and generated molecules (MST). We compute the Morgan fingerprints (Zhong & Guan, 2023) of all the generated and test dataset molecular graphs and determine the maximum Tanimoto similarity between each generated molecule and the test set.
@@
To assess the quality of the generated 3D molecular conformations, we employ the Wasserstein metrics introduced in (Vignac et al., 2023b), and compute bond angle (BA) and bond length (BL) W1 distance between generated and data distributions. We additionally compute conformational energies (CEMMFF and CExTB) W1 distances, based on the MMFF and GFN2-xTB (Bannwarth et al., 2019) energies, respectively, of generated and data conformations. Since point-cloud methods may create large outliers and W1 distances are outlier sensitive, we restrict generated angles and lengths to those generated within the data limits (see Appendix H).
@@ 6 Experiments and results
We report the basic molecule properties and molecule graph quality metrics in Tables 1 and 2, while molecule conformation and conditional generation results can be found in Appendix A.1 and A.2.
Previous methods have employed many different settings for molecule generation. While all of them may be valid approaches, it makes benchmarking challenging. We note in Appendix I how explicit aromatic generation makes it difficult to compute and compare certain metric values against Kekule representations (used in this work). In Tables 1 and 2, we annotate explicit aromatic formulations with ()($\ddagger$)( ‡ ). Methods that employ post-generation corrections or algorithmic bond extractions are annotated with ()($\dagger$)( † ). The methods we tested ourselves are marked with (*).
@@ 6.1 QM9
QM9 (Ramakrishnan et al., 2014) is a small molecule dataset, containing 134k molecules of at most 29 atoms (including hydrogen). We use the splits from (Hoogeboom et al., 2022), with 100k, 18k, and 10k molecules for training, validation, and testing respectively. We use 98522 molecules out of 100k, selected s.t. $\mathbf{m}_{i}\pm 2\sigma$ of all atom RBF components fit within the evaluated field positions $\mathcal{G}$.
Basic molecular properties are reported in Table 1. We achieve state-of-the-art performance in terms of molecule and atom neutrality. In Table 2, the maximum test similarity and the bond distribution are competitive to the point-cloud approaches, while MiDi performs very well on TVa.
This is likely an effect of our model trading the correct atom count generation for improved neutrality. Point cloud methods are hard-constrained to generate specific numbers of atoms, while our method may remove or add atoms in favor of other properties. In Appendix G, we report a small shift in the generated atom count distribution compared to the desired conditioning, while increasing the strength of classification guidance $\beta$ in Table 3 significantly decreases TVa.
Although our NN architecture is not invariant to SE(3) transformations, our method achieves satisfactory conformational accuracy, with a very low distance between generated and data distributions. We report the W1 distances for bond angles, bond distances, and conformation energy distributions in Appendix A.1, Table 4, and show the corresponding CDFs in Figures 4 and 5.
Table 1: Basic molecule properties for molecules generated by models trained on the QM9 and GEOM-Drugs datasets, based on 10k samples. We report the mean and standard deviation of the metrics for three different runs (each generating 10k samples) for QM9. The remark column (R) symbols represent: (*) re-tested by us ; ($\dagger$) post-processing corrections or algorithmic bond extraction; ()($\ddagger$)( ‡ ) explicit aromatic bonds. Results for methods with $\dagger$ or $\ddagger$ are in italics since they cannot be directly compared with purely generative models. An H column $\checkmark$ denotes explicit hydrogen generation.
QM9 GEOM-Drugs
Neutrality Neutrality
Model H R Atom(%) Mol(%) Val(%) R Atom(%) Mol(%) Val(%)
MolDiff ($\dagger$, $\ddagger$) - - 97.0 - - - -
VoxMol ($\dagger$) 99.2 89.3 98.7 ($\dagger$) 98.1 75.0 93.4
MiDi (*) 98.3 84.5 96.7 (*,$\ddagger$)‡ ) 96.8 31.3 62.7
JODO - 98.5 88.2 98.1 ($\ddagger$) 97.0 36.7 77.8
EDM (*) 98.1 81.9 91.6 (*) 81.0 0.5 3.0
GDM-AUG - 97.6 71.6 90.4 - - -
EDM-Bridge - 98.8 84.6 92.0 - - - -
G-Schnet - 95.7 68.1 85.5 - - - -
GeoLDM - 98.9 89.4 93.8 - 84.4 - 99.3
FMG - 99.4$\pm$ 0.1 91.5$\pm$ 1.4 98.8 $\pm$ 0.3 - 94.7 16.5 64.9
Data - 99.9 99.6 99.6 - 99.4 86.2 89.6
MolDiff - - - - (,)(*$\ddagger$)( ∗ , ‡ ) - - 85.0
FMG - - - - - - - 80.0
Table 2: Molecule graph quality of generated samples from models trained on the QM9 and GEOM datasets, based on 10k samples. We report the mean and standard deviation of the metrics for three different runs (each generating 10k samples) for QM9. See Table 1 for description of symbols.
QM9 GEOM-Drugs
Model H R MST (↑) TVa(↓) TVb(↓) R MST (↑) TVa(↓) TVb(↓)
MiDi (*) 0.57 0.19 0.81 (*$\ddagger$) 0.53 1.07 -
EDM (*) 0.55 0.49 0.98 (*) 0.46 1.63 5.72
JODO (*) 0.58 0.19 0.89 (*, $\ddagger$) 0.58 1.07 -
FMG - 0.54$\pm$ 0.00 0.52 $\pm$ 0.02 0.95 $\pm$ 0.01 - 0.49 2.4 4.78
MolDiff - - - - (*$\ddagger$) 0.62 2.03 -
FMG - - - - 0.50 0.63 4.58
6.2 QM9 conditional generation
We employ the same setting as in previous work (Hoogeboom et al., 2022), training molecule property discriminator models on 50% of the QM9 data, and conditional generation models on the remaining 50%. Our generated molecules are then tested using the classifiers. Further details of the experiment and the results are reported in Appendix A.2.
We compare the difference between the highest and lowest occupied molecular energies $\Delta\epsilon$ and the polarizability $\alpha$ of the generated molecules, compared to the true, conditional values. FMG outperforms EDM in terms of molecular stability, while EDM is better in its fidelity to the conditioning variables $c$.
Additionally, our method generates similar fidelity to the conditioning variable $c$, even without classification-free guidance on the number of atoms $N_{a}\sim p(N_{a}|c)$. This could prove useful in situations where the conditional $p(N_{a}|c)$, cannot be accurately determined based on data statistics (e.g., for generating ligands that interact with specific proteins).
6.3 GEOM-Drugs
The Geometric Ensemble Of Molecules (GEOM) dataset (Axelrod & Gómez-Bombarelli, 2022) contains molecules of up to 181 atoms and 37 million conformations along with their corresponding energies. Following previous work (Hoogeboom et al., 2022; Peng et al., 2023), we select the 30 lowest energy conformers for each molecule graph. Similar to (Peng et al., 2023), we also restrict the data points to only those molecules that contain atoms from {H, C, N, O, F, P, S, Cl}. We extract those molecules that can be contained in the pre-defined evaluated positions that determine the tensors $\mathbf{u}_{a}$ and $\mathbf{u}_{b}$, similar to our QM9 experiments. After filtering, the data contains 96.7% of the training molecular conformations and 99.7% unique molecular graph coverage.
We report basic molecule properties in Table 1, where our method achieves highly competitive neutral atoms and molecule percentages. Because of post-generation bond extractions and corrections (marked with ($\dagger$)) and other bond representation choices, it remains unfortunately difficult to reliably compare against all methods (please refer to Appendix I for details). In Table 2, we see the same effect mentioned in the QM9 experiments, where the method has a worse TVa metric, likely caused by our method favoring other metrics to correct atom counts. When H atoms are not explicitly generated, our generated atom distribution becomes significantly closer to the data.
The conformational properties of the generated molecules are detailed in Appendix A.1. Our method performs well, even with significantly more complex molecules.
6.4 E(3) invariance and chirality analysis
In this section, we empirically confirm that E(3) invariant implementations lead to invariance wrt enantiomers and show that our proposed method captures all molecular geometries. Our setting consists of a synthetic dataset of 48 chiral molecules with a skewed enantiomer distribution. Details of the experiment are provided in Appendix B.2.
Figure 3 shows the KDEs of generated tetrahedron determinants of ${\bm{R}}=[\mathbf{m}_{\texttt{O}}-\mathbf{m}_{\texttt{C}},\mathbf{m}_{\texttt{N}}-\mathbf{m}_{\texttt{C}},\mathbf{m}_{\texttt{H}}-\mathbf{m}_{\texttt{C}}]^{T}\in\mathbb{R}^{3\times 3}$. The E(3) equivariant generative method (EDM) has $p_{\phi}(\det{\bm{R}})=p_{\phi}(-\det{\bm{R}})$, which empirically attests Proposition 1.
Figure 3: Kernel density estimation of $\det{\bm{R}}$, resulted from the data and generated molecules. EDM is unable to distinguish and correctly generate the enantiomer distribution, while the E(3) variant method (GNN) generates improper, 0 volume conformations. FMG matches the enantiomer distribution.
We also test a point-cloud method that is variant to reflections. The work of Hoogeboom et al. (2022) (GDM-AUG in Figure 3) provides a variant of their method, where atom positions $\{\mathbf{m}_{i}\}$ are concatenated to node features. Interestingly, although this method has $p_{\phi}(\det{\bm{R}})\neq p_{\phi}(-\det{\bm{R}})$, it generates many unnatural conformations, with tetrahedrons of approximately 0 volume.
Finally, FMG captures the distribution well.
6.5 Ablation
We determine the aspects of our method that contribute the most to our competitive performance. In Table 3, FMGaug receives augmented fields based on randomly rotated sets of atom coordinates $\{\mathbf{m}_{i}\}$ during training, FMGβ, $\beta\in\{0,2,5\}$ varies the degree of classification free guidance on the number of atoms (see Section 4.4), and FMGdist is the performance of the molecular graphs, where bonds are extracted according to bond pairwise atom distances, without considering $\mathbf{u}_{b}$ channels. The rotation alignment is the main factor that increases the neutrality metric to state-of-the-art performance, as FMGaug performance drops significantly, while classification guidance improves TVa.
For FMGdist, we determine bonds based on the lookup table from (Hoogeboom et al., 2022). This confirms that our molecular conformations analyzed in Appendix A.1 are remarkably good since the model’s neutrality remains SOTA even when bonds are solely based on the atom positions $\{\mathbf{m}_{i}\}$.
Table 3: FMG performance with various design choices.